# Fast empirical scenarios

### v2

@@ Abstract.
We seek to extract a small number of representative scenarios from large panel data that are consistent with sample moments. Among two novel algorithms, the first identifies scenarios that have not been observed before, and comes with a scenario-based representation of covariance matrices. The second proposal selects important data points from states of the world that have already realized, and are consistent with higher-order sample moment information. Both algorithms are efficient to compute and lend themselves to consistent scenario-based modeling and multi-dimensional numerical integration. Extensive numerical benchmarking studies and an application in portfolio optimization favor the proposed algorithms.
@@ 1. Introduction
Multi-dimensional data in various fields require efficient processing for informed decision making. In many practical applications, the moments of the sample realizations induced by the sample distribution, together with the sample realizations themselves, become central to inference tasks. For instance, portfolio optimization involving risky assets necessitates covariance matrices or even higher-order and/or non-linear functions of the moments, while the corresponding risk management is often scenario-based. The factor structure in interest rates also suggests a scenario-based approach, see Engle et al. (2017). More generally, the topic of summarizing information by replacing large samples of data with a small number of carefully weighted scenarios have found traction in many different communities, eg. scalable Bayesian statistics, see Huggins et al. (2016), clustering and optimization, see Feldman et al. (2020), etc.
@@ 2.1. Background
Let $\Omega\subset\mathbb{R}^{d}$ and $\mathscr{B}$ denote the Borel $\sigma$-algebra on $\Omega$. Suppose that we are given a finite real sequence $\boldsymbol{y}=\big{(}y_{\boldsymbol{\alpha}}\big{)}_{|\boldsymbol{\alpha}|\leq q}$ indexed by $\boldsymbol{\alpha}\mathrel{\mathrel{\mathop{:}}=}(\alpha_{1},\ldots,\alpha_{d})\in\mathbb{N}^{d}$ with $|\boldsymbol{\alpha}|\mathrel{\mathrel{\mathop{:}}=}\alpha_{1}+\cdots+\alpha_{d}\leq q$, where $q\in\mathbb{N}$. We say that $\boldsymbol{y}$ has a representing measure if there exists a Borel measure $\mu:\mathscr{B}\to[0,\infty)$ such that
@@ 2.3. Covariance scenarios
output: scenarios $\Xi\mathrel{\mathrel{\mathop{:}}=}\big{\{}\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{r}\big{\}}$
@@ 3.2. Reformulation in RKHS
For completeness, we provide the definition of the RKHS below, see Paulsen (2016), Berlinet and Thomas-Agnan (2011), for further details.
@@
Let $\big{(}\mathcal{H},\langle{\cdot},{\cdot}\rangle_{\mathcal{H}}\big{)}$ be a Hilbert space of real-valued functions on a non-empty set $\Omega$. A function $\mathcal{K}:\Omega\times\Omega\longrightarrow\mathbb{R}$ is a reproducing kernel of the Hilbert space $\mathcal{H}$ iff
@@
The function $\mathcal{K}:X\times X\to\mathbb{R}$ as defined in (3.7) is a symmetric and positive type function. Moreover, $\mathcal{K}$ has the reproducing property on $X$,
@@
[…]
@@ 4. Algorithms for scenario representation
In this section, we present the algorithms for the two stages as mentioned in Section 3. The first algorithm, orthogonal matching pursuit (OMP), see Pazouki and Schaback (2011), is devised to extract the scenarios in a greedy manner. The second is the alternating direction method of multipliers (ADMM), see Boyd et al. (2011), and is used for retrieving the corresponding probabilities.
@@ 4.1. Step 1: Scenario extraction
output: index set $\pi$ corresponding to the sparse scenarios,
@@
$\boldsymbol{\ell}_{m}\mathrel{\mathrel{\mathop{:}}=}\frac{1}{\sqrt{d_{\pi(m)}}}\Big{(}{\boldsymbol{K}}-{\boldsymbol{L}_{m-1}}{\boldsymbol{L}}^{\top}_{m-1}\Big{)}\boldsymbol{e}_{\pi(m)}\quad\ \text{and}\quad\boldsymbol{b}_{m}\mathrel{\mathrel{\mathop{:}}=}\frac{1}{\sqrt{d_{\pi(m)}}}\Big{(}{\boldsymbol{I}}-{\boldsymbol{B}_{m-1}}{\boldsymbol{L}}^{\top}_{m-1}\Big{)}\boldsymbol{e}_{\pi(m)}$
@@ 4.2. Step 2: Retrieval of probability weights
Having extracted the set of $m$ scenarios $\{\boldsymbol{\xi}_{1},\ldots,\boldsymbol{\xi}_{m}\}=\{\boldsymbol{x}_{\pi(1)},\ldots,\boldsymbol{x}_{\pi(m)}\}$, we proceed to enforce the probability simplex constraints. We deploy the ADMM algorithm, see Boyd et al. (2011); Parikh and Boyd (2014). This approach can be used for projecting least-squares solutions of linear systems onto convex sets. Denote the set of the two simplex constraints as
@@
output: $m$ scenarios $\Xi\mathrel{\mathrel{\mathop{:}}=}\big{\{}\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{m}\big{\}}$ and their probabilities $\{\lambda_{1},\ldots,\lambda_{m}\}$
@@
We define the two respective cost functions $g_{\lambda},h_{\lambda}:\mathbb{R}^{N}\longrightarrow\mathbb{R}$ as
@@
Note that in the above proof, we do not assume any positivity of the weight vector $\boldsymbol{w}$. The result therefore has far-reaching consequences for practical purposes. Foremostly, it shows that it is not possible to obtain optimal quadrature rules (with few atoms) in general that match the empirical moments of polynomials. This fact severely restricts the pool of candidate algorithms that can be deployed for efficient recovery of a few scenarios from the empirical moments. In particular, many popular first-order proximal gradient methods, like the FISTA, see Beck and Teboulle (2009); Nesterov (1983), POGM, see Taylor et al. (2017); Donghwan and Fessler (2018), etc. to name a few cannot be used in this context.
@@ 5.1. Multivariate Gaussian mixture distributions
We first examine the different algorithms on simulated data sampled from a Gaussian mixture model with different numbers of dimensions $d$ and having different number of clusters $c$, components of the mixture distribution, that induce multiple modes into the joint distribution. We test for $d=2,5,10,20,50,100$ and $c=5,10,20,25,50,100,150,200$. To investigate the efficacy of the algorithms, we do not use a higher $c$ for relatively lower $d$ and vice-versa. The means of the different clusters are taken to be random vectors of uniformly distributed numbers in the interval $(-50,50)$. The variance-covariance matrices for the different clusters are either (i) randomly generated positive definite matrices with $\mathcal{N}(1,1)$-distributed eigenvalues, or (ii) the identity matrix. Except in the case of unit variance-covariance, the different clusters have different variance-covariance matrices. The mixing proportions for the different clusters are taken to be either (i) random or, (ii) equal. We can categorize them as
@@
For each of the above cases, $100$ data sets are randomly constructed, each containing $10000$ samples. For each data set containing the samples, we calculate the sample moments $\widehat{\boldsymbol{y}}$ up to order $2q$ with $q=1,2$ which give rise to the empirical moment matrices $\boldsymbol{M}_{\widehat{\boldsymbol{y}}}$ of orders $1$ and $2$ respectively. The $100$-dimensional and $50$-dimensional data sets are computed for $q=1$ only, still resulting in $5151$-dimensional, and $1326$-dimensional moment matrices, respectively. The biggest moment matrix for $q=2$ is computed for twenty dimensions, resulting in a $10262$-dimensional moment matrix. It is noteworthy to point out that there is, theoretically speaking, no bound on the degree to which we can calculate the moments. However, the computational cost of obtaining higher-order moment information is high.
@@ 5.1.3. Computation time
Finally, we test how fast the algorithms are in computing the scenarios from large empirical data sets, keeping practical applications in mind. To that end, we plot the CPU run times of the different algorithms against the different dimensions and number of clusters for $q=1,2$ in Figure 5 and Figure 6 respectively. We take the $y$-axis to be the computation times in the log scale. The $x$-axis contains the different dimensions which is a categorical variable. The color bar depicts the different number of clusters and is also taken as a categorical variable.
@@ 5.2. Portfolio optimization with CVaR constraints
In this section, we discuss an application of the proposed OMP algorithm in finance using portfolio excess return data from Fama-French. The data set contains more than $25000$ daily excess returns of $25$ portfolios. An important aspect of portfolio risk management is to address extreme outcomes, wherein investors are concerned by the multivariate nature of risk and the scale of the portfolios under consideration. For decision-making, any sensible strategy involves dimension reduction and modeling the key features of the overall risk landscape. To this end, we consider here the problem of maximizing expected portfolio returns with expected shortfall constraints as a proxy for the entailing risk. Expected shortfall, also known as conditional value-at-risk (CVaR), is a coherent risk measure that quantifies the tail risk an investment portfolio has, see McNeil et al. (2015). Precisely, we consider the following optimization problem:
(5.2) $\begin{split}\boldsymbol{w}^{*}&=\underset{\boldsymbol{w}\in\mathbb{R}^{d}}{\operatorname{argmin}}\quad-\boldsymbol{w}^{\top}\boldsymbol{\mu}\\ \text{subject to: }&\text{CVaR}_{\alpha}\leq\delta,\quad\boldsymbol{w}\in W.\end{split}$
@@
We use the reformulation of the CVaR as in Uryasev and Rockafellar (2001) to find a global solution. While generally for the optimization approach to work, simulations using a Monte-Carlo approach or bootstrapping of historical data are required, we use our generated scenarios, using the OMP, to showcase the applicability in the case of non-smooth optimization as well.
Figure 7. Daily mean returns versus expected shortfall at $95\%$ confidence level for $1\%,2\%,5\%$ quantile levels of the naive portfolio rule. The data are daily portfolio returns from the $25$-dimensional Fama-French portfolios.
To that end, we split the portfolio data set into a training set of $10000$ observations and a testing set of the rest of the observations. We first extract the scenarios and the corresponding probabilities from the training data set using the OMP algorithm, see Algorithm 4.1. We then solve problem (5.2) using the scenarios and the approach suggested in Uryasev and Rockafellar (2001), taking the conditional value-at-risk constraint to be at $95\%$ confidence level, the $\text{CVaR}_{\alpha}$ for $\alpha=0.95,0.98,0.99$ for the returns of the naive portfolio rule ${\boldsymbol{w}}=(1/d){\boldsymbol{1}}$. We perform back-testing using the optimized portfolio weights on the test sample observations. We perform $100$ simulations and observe the distribution of the expected daily returns versus the CVaR or the expected shortfall for both the training and testing samples in Figure 7. All the observations are considered as percentages, the color bar indicates the $\alpha$ level considered for the optimization problem. The figure shows that the optimized portfolios perform well on the testing set as well, achieving a yearly mean return of about $14\%$ to $17\%$ in general; hence, our realized scenarios can still capture the behavior of the underlying asset returns considerably well, even when trying to match moments of non-smooth functions of polynomials.

### v3

@@ Abstract.
We seek to extract a small number of representative scenarios from large panel data that are consistent with sample moments. Among two novel algorithms, the first identifies scenarios that have not been observed before, and comes with a scenario-based representation of covariance matrices. The second proposal selects important data points from states of the world that have already realized, and are consistent with higher-order sample moment information. Both algorithms are efficient to compute and lend themselves to consistent scenario-based modeling and multi-dimensional numerical integration that can be used for interpretable decision-making under uncertainty. Extensive numerical benchmarking studies and an application in portfolio optimization favor the proposed algorithms.
@@ 1. Introduction
Multi-dimensional data in various fields require efficient processing for informed decision-making. In many practical applications, the moments of the sample realizations induced by the sample distribution, together with the sample realizations themselves, become central to inference tasks. For example, investors care not only about the variance of outcomes but also about the higher-order moments of the distribution of outcomes, in particular, about the possibility of extremely adverse outcomes. Similarly, in the case of estimating the stochastic discount factor (SDF) for asset pricing, one must ensure that it incorporates adequate information about the higher moments of the return distribution as it is important for modeling tail risk, see Schneider and Trojani (2015); Almeida and Garcia (2017); Almeida et al. (2022). As a result, portfolio optimization involving risky assets necessitates covariance matrices or even higher-order and/or non-linear functions of the moments, while the corresponding risk management is often scenario-based. The factor structure in interest rates also suggests a scenario-based approach, see Engle et al. (2017). More generally, the topic of summarizing information by replacing large samples of data with a small number of carefully weighted scenarios has found traction in many different communities, eg.  scalable Bayesian statistics, see Huggins et al. (2016), clustering and optimization, see Feldman et al. (2020), etc. In the discipline of explainable artificial intelligence, the scenarios’ probabilities can be used to define observation-specific explanations that give rise to a novel class of surrogate models, see Ghidini et al. (2024).
@@ 2.1. Background
Let $\Omega\subset\mathbb{R}^{d}$ and $\mathscr{B}$ denote the Borel $\sigma$-algebra on $\Omega$. Suppose that we are given a finite real sequence $\boldsymbol{y}=\big{(}y_{\boldsymbol{\alpha}}\big{)}_{|\boldsymbol{\alpha}|\leq q}$ indexed by $\boldsymbol{\alpha}\mathrel{\mathrel{\mathop{:}}=}(\alpha_{1},\ldots,\alpha_{d})\in\mathbb{N}^{d}$ with $|\boldsymbol{\alpha}|\mathrel{\mathrel{\mathop{:}}=}\alpha_{1}+\cdots+\alpha_{d}\leq q$, where $q\in\mathbb{N}$. We say that $\boldsymbol{y}$ has a representing measure if there exists a Borel measure $\mu\colon\mathscr{B}\to[0,\infty)$ such that
@@ 2.3. Covariance scenarios
output: scenarios $\Xi\mathrel{\mathrel{\mathop{:}}=}\big{\{}\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{r}\big{\}}\subset\mathbb{R}^{d}$
@@ 3.2. Reformulation in RKHS
For completeness, we provide the definition of the RKHS below, see Paulsen and Raghupathi (2016), Berlinet and Thomas-Agnan (2011), for further details.
@@
Let $\big{(}\mathcal{H},\langle{\cdot},{\cdot}\rangle_{\mathcal{H}}\big{)}$ be a Hilbert space of real-valued functions on a non-empty set $\Omega$. A function $\mathcal{K}\colon\Omega\times\Omega\longrightarrow\mathbb{R}$ is a reproducing kernel of the Hilbert space $\mathcal{H}$ iff
@@
The function $\mathcal{K}\colon X\times X\to\mathbb{R}$ as defined in (3.7) is a symmetric and positive type function. Moreover, $\mathcal{K}$ has the reproducing property on $X$,
@@
$\frac{\sigma_{\text{min}}(\boldsymbol{M}^{\dagger})}{N}\cdot\big{\|}{\boldsymbol{V}}^{\top}{\boldsymbol{w}}-\widehat{\boldsymbol{y}}\big{\|}_{2}^{2}\hskip 3.0pt\leq\hskip 3.0pt\big{\|}\boldsymbol{Q}^{\top}\boldsymbol{w}-\widetilde{\boldsymbol{y}}\big{\|}_{2}^{2}\leq\frac{\sigma_{\text{max}}(\boldsymbol{M}^{\dagger})}{N}\cdot\big{\|}{\boldsymbol{V}}^{\top}{\boldsymbol{w}}-\widehat{\boldsymbol{y}}\big{\|}_{2}^{2}.$
@@ 4. Algorithms for scenario representation
In this section, we present the algorithms for the two stages as mentioned in Section 3. The first algorithm, orthogonal matching pursuit (OMP), see Pazouki and Schaback (2011), is devised to extract the scenarios in a greedy manner. The second is the alternating direction method of multipliers (ADMM), see Boyd (2011), and is used for retrieving the corresponding probabilities.
@@ 4.1. Step 1: Scenario extraction
output: index set $\pi$ corresponding to the sparse scenarios $\Xi\mathrel{\mathrel{\mathop{:}}=}\{\boldsymbol{\xi}_{1},\ldots,\boldsymbol{\xi}_{m}\}\subset\mathbb{R}^{d}$,
@@
$\boldsymbol{\ell}_{m}\mathrel{\mathrel{\mathop{:}}=}\frac{1}{\sqrt{d_{m-1,\pi(m)}}}\Big{(}{\boldsymbol{K}}-{\boldsymbol{L}_{m-1}}{\boldsymbol{L}}^{\top}_{m-1}\Big{)}\boldsymbol{e}_{\pi(m)}\quad\ \text{and}\quad\boldsymbol{b}_{m}\mathrel{\mathrel{\mathop{:}}=}\frac{1}{\sqrt{d_{m-1,\pi(m)}}}\Big{(}{\boldsymbol{I}}-{\boldsymbol{B}_{m-1}}{\boldsymbol{L}}^{\top}_{m-1}\Big{)}\boldsymbol{e}_{\pi(m)}$
@@ 4.2. Step 2: Retrieval of probability weights
Having extracted the set of $m$ scenarios $\Xi\mathrel{\mathrel{\mathop{:}}=}\{\boldsymbol{\xi}_{1},\ldots,\boldsymbol{\xi}_{m}\}=\{\boldsymbol{x}_{\pi(1)},\ldots,\boldsymbol{x}_{\pi(m)}\}$, we proceed to enforce the probability simplex constraints. We deploy the ADMM algorithm, see Boyd (2011); Parikh (2014). This approach can be used for projecting least-squares solutions of linear systems onto convex sets. Denote the set of the two simplex constraints as
@@
output: $m$ scenarios $\Xi\mathrel{\mathrel{\mathop{:}}=}\big{\{}\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{m}\big{\}}\subset\mathbb{R}^{d}$ and their probabilities $\{\lambda_{1},\ldots,\lambda_{m}\}\subset\Delta$
@@
We define the two respective cost functions $g_{\lambda},h_{\lambda}\colon\mathbb{R}^{N}\longrightarrow\mathbb{R}$ as
@@
Note that in the above proof, we do not assume any positivity of the weight vector $\boldsymbol{w}$. The result therefore has far-reaching consequences for practical purposes. Foremostly, it shows that it is not possible to obtain optimal quadrature rules (with few atoms) in general that match the empirical moments of polynomials. This fact severely restricts the pool of candidate algorithms that can be deployed for efficient recovery of a few scenarios from the empirical moments. In particular, many popular first-order proximal gradient methods, like the FISTA, see Beck and Teboulle (2009); Nesterov (1983), POGM, see Taylor et al. (2017); Kim and Fessler (2018), etc. to name a few cannot be used in this context.
@@ 5.1. Multivariate Gaussian mixture distributions
We first examine the different algorithms on simulated data sampled from a Gaussian mixture model with different numbers of dimensions $d$ and having different numbers of clusters $c$, components of the mixture distribution, that induce multiple modes into the joint distribution.
Figure 1. The OMP scenarios (red dots) from $10000$ random samples of a bivariate Gaussian mixture distribution with their contour lines. The PDF of the distribution is given in the first tile.
We test for $d=2,5,10,20,50,100$ and $c=5,10,20,25,50,100,150,200$. To investigate the efficacy of the algorithms, we do not use a higher $c$ for relatively lower $d$ and vice-versa. The means of the different clusters are taken to be random vectors of uniformly distributed numbers in the interval $(-50,50)$. The variance-covariance matrices for the different clusters are either (i) randomly generated positive definite matrices with $\mathcal{N}(1,1)$-distributed eigenvalues, or (ii) the identity matrix. Except in the case of unit variance-covariance, the different clusters have different variance-covariance matrices. The mixing proportions for the different clusters are taken to be either (i) random or, (ii) equal. We can categorize them as
@@
For each of the above cases, $100$ data sets are randomly constructed, each containing $10000$ samples. For each data set containing the samples, we calculate the sample moments $\widehat{\boldsymbol{y}}$ up to order $2q$ with $q=1,2$ which give rise to the empirical moment matrices $\boldsymbol{M}_{\widehat{\boldsymbol{y}}}$ of orders $1$ and $2$ respectively. The $100$-dimensional and $50$-dimensional data sets are computed for $q=1$ only, still resulting in $5151$-dimensional, and $1326$-dimensional moment matrices, respectively. The biggest moment matrix for $q=2$ is computed for twenty dimensions, resulting in a $10262$-dimensional moment matrix. It is noteworthy to point out that there is, theoretically speaking, no bound on the degree to which we can calculate the moments. However, the computational cost of obtaining higher-order moment information is high. We provide a visualization for the case $d=2$ where we first generate $10000$ random samples from a bivariate Gaussian mixture distribution with $20$ clusters, and then retrieve the OMP scenarios as in Figure 1.
@@ 5.1.3. Computation time
Finally, we test how fast the algorithms are in computing the scenarios from large empirical data sets, keeping practical applications in mind. To that end, we plot the CPU run times of the different algorithms against the different dimensions and number of clusters for $q=1,2$ in Figure 6 and Figure 7 respectively. We take the $y$-axis to be the computation times in the log scale. The $x$-axis contains the different dimensions, which is a categorical variable. The color bar depicts the different number of clusters and is also taken as a categorical variable.
@@ 5.2. Portfolio optimization with CVaR constraints
In this section, we discuss an application of the proposed OMP algorithm in finance using excess return data from Fama-French. The data set contains more than $25000$ daily excess returns of $25$ financial assets. A portfolio, or trading strategy, is defined through the number and choice of assets. An important aspect of portfolio risk management is to address extreme outcomes, wherein investors are concerned by the multivariate nature of risk and the scale of the portfolios under consideration. For decision-making, any sensible strategy involves dimension reduction and modeling the key features of the overall risk landscape. To this end, we consider here the problem of maximizing expected portfolio returns with expected shortfall constraints as a proxy for the entailing risk. Expected shortfall, also known as conditional value-at-risk (CVaR), is a coherent risk measure that quantifies the tail risk an investment portfolio has, see McNeil et al. (2015). In 2016, the Basel committee hosted by the Bank of International Settlements (BIS) proposed an internationally agreed set of regulatory measures in response to the financial crisis of 2007-2009; one of them was that the market risk capital of banks should be calculated with CVaR or expected shortfall, see Basel Committee on Banking Supervision (2019) (BCBS). Nevertheless, methods for portfolio optimization that rely on such a downside risk measure as CVaR, are often difficult to implement, do not scale efficiently, and may result in less-than-ideal portfolio allocations. To resolve these issues and construct an optimal portfolio that minimizes shortfall risk, we consider the following optimization problem:
(5.2) $\begin{split}\boldsymbol{w}^{*}&=\underset{\boldsymbol{w}\in\mathbb{R}^{d}}{\operatorname{argmin}}\quad-\boldsymbol{w}^{\top}\boldsymbol{\mu}\\ \text{subject to: }&\text{CVaR}_{\alpha}\leq\delta,\quad\boldsymbol{w}\in W,\end{split}$
@@
We use the reformulation of the CVaR as in Uryasev and Rockafellar (2001) to find a global solution. While, generally for the optimization approach to work, simulations using a Monte-Carlo approach or bootstrapping of historical data are required, we use our generated scenarios, using the OMP, to showcase the applicability in the case of non-smooth optimization as well. The scenarios from the OMP capture the underlying tail risk, which leads to efficient portfolios whose risk lies below a certain threshold.
Figure 8. Daily mean returns versus CVaR for the worst $1\%,2.5\%$ and $5\%$ cases of the expected portfolio loss applied to the training and testing samples with both optimized portfolios and the naive portfolio. The data are daily returns from the $25$-dimensional Fama-French portfolios.
To that end, we split the portfolio data set into a training set of $10000$ observations and a testing set of the rest of the observations. We first extract the scenarios and the corresponding probabilities from the training data set using the OMP algorithm, see Algorithm 4.1. We then solve problem (5.2) using the scenarios and the approach suggested in Uryasev and Rockafellar (2001), taking $\delta$ to be the conditional value-at-risk for the the naive portfolio rule ${\boldsymbol{w}}=(1/d){\boldsymbol{1}}$ at the confidence levels $\alpha=0.95,0.975,0.99$, which correspond to the expected loss in the worst $5\%,2.5\%$ and $1\%$ cases respectively. The $\operatorname{CVaR}_{\alpha}$ in the constraint is the corresponding conditional value-at-risk of the expected portfolio loss in the worst $5\%,2.5\%$ and $1\%$ case as well. We perform back-testing using the optimized portfolio weights on the test sample observations. We perform $100$ simulations and observe the distribution of the expected daily returns versus the CVaR or the expected shortfall for both the training and testing samples in Figure 8. Furthermore, we plot the same for the case of the equally weighted portfolios. All the observations are considered as percentages, the color bar indicates the $\alpha$ level considered for the optimization problem. The figure shows that the optimized portfolios perform well on the testing set as well, achieving a yearly mean return of about $12\%$ to $14\%$ in general; hence, our realized scenarios can still capture the behavior of the underlying asset returns considerably well, even when trying to match moments of non-smooth functions of polynomials.